In a recent study, 30% of people surveyed indicate chocolate was their favorite flavor of ice cream. Suppose we select a sample of 9 people and ask them to name their favorite flavor of ice cream.

  1. How many of those in the sample would you expect to name chocolate? (Round your answer to 1 decimal place.)

  2. What is the probability exactly three of those in the sample name chocolate? (Round the probability to 5 decimal places and the final answer to 4 decimal places.)

  3. What is the probability three or more name chocolate? (Round your probability to 4 decimal places.)
Answer #1

Ans:

Use binomial distribution with n=9 and p=0.30

a)

Expected number who name chocolate=np=9*0.30=2.7

b)

P(x=3)=9C3*0.3^3*(1-0.3)^6=0.2668

c)

P(3 or more)=P(x>=3)=1-P(x<=2)

=1-P(x=0)-P(x=1)-P(x=2)

=1-(1-0.3)^9-9C1*0.3*(1-0.3)^8-9C2*0.3^2*(1-0.3)^7

= 0.5372